In a statistic examination, the mean was 55 and the standard deviation was 10. Determine the standard scores of two students whose grade were 75 and 42 respectively.


1
Expert's answer
2021-06-21T11:19:30-0400
"z=\dfrac{x-\mu}{\sigma}"

Given "\mu=55, \sigma=10"


"x=75: z=\dfrac{75-55}{10}=2"

"x=42: z=\dfrac{42-55}{10}=-1.3"